OpenStudy (anonymous):

Please Help!! I've tried numerous times and am still getting wrong answer....thanks :) Use left and right endpoints and the given number of rectangles to find two approximations of the area of the region between the graph of the function and the x-axis over the given interval. g(x) = 2x^2 − x − 1, [3, 5], 4 rectangles

OpenStudy (perl):

i can help , bear with me, helping other people

OpenStudy (anonymous):

ok, thanks,i have the leftend point...but not the right end point....i don't know where i went wrong there

OpenStudy (perl):

i am here, but i am helping multiple people, can you wait

OpenStudy (perl):

just keep tagging me so i dont forget you

OpenStudy (anonymous):

yes, i can

OpenStudy (anonymous):

ok

OpenStudy (perl):

ok im back

OpenStudy (perl):

delta x = (b-a)/n delta x = ( 5 - 3 ) / 4 = 1/2

OpenStudy (anonymous):

yeah, i got that

OpenStudy (perl):

so first lets look at the x axis we have x = 3, 3.5, 4, 4.5, 5 (note that there are 5 values here)

OpenStudy (anonymous):

ok...i have that part and to find y you plug that into the function, correct?

OpenStudy (perl):

Left endpoint sum = 1/2 * f(3) + 1/2 * f(3.5) + 1/2 * f(4) + 1/2 * f(4.5) Right endpoint sum = 1/2 * f(3.5) + 1/2 * f(4) + 1/2 * f(4.5) +1/2 * f(5)

OpenStudy (anonymous):

for the right, i got 63 and thats correct...but for the left i got 48.5 and thats wrong @perl

OpenStudy (perl):

ok checking

OpenStudy (perl):

i would also factor out 1/2

OpenStudy (anonymous):

ok..i'll try that

OpenStudy (perl):

also ill make a little table

OpenStudy (perl):